A cube has a volume of 7 to the 3rd power cubic centimeters. What is it's surface area?
Answer by addingup(3677)

This cube has 6 faces, and all side lengths are equal.
:
The volume of a cube with sides of length L is given by L^3.
:
The problem says that the volume is 7^3, which means that each side is 7
:
The surface area of a cube of length L is 6*L^2. Since L = 7, the surface area is:
6*7^2 = 6*49 = 294 cm^2